Project

General

Profile

Actions

Commissioning and Ops Scheduling meeting : 15/12/2015

To be held in CR10 in the ATLAS Building @ 15:30 pm

Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402

Agenda:

  • Review of tasks and activities identified at last meeting
    • Vacuum / Hydrogen system work : Mark Tucker
    • Running with Xe in the absorber : MOM
  • Upcoming activities : chaired by DC
    • This week's MICE Hall work plan (Craig)
    • Evacuating absorber to provide a baseline scattering measurement : do we and when?
    • Activities after running finishes on Friday 18th
      • test of the Q4 flush - take magnet to high current and see if it trips.
      • Temperature readout investigation/swap of Q4 / Q5 sensors (Henry)
      • 50A test on the DS (Mike)
    • Network reconfiguration for C&M. It is estimated that this may take up to a week of work. When do we schedule?
    • Christmas shutdown schedule and plans : when is "end of play"?
      • What needs to be made safe before leaving for Christmas?
  • AOB
    • To-do list for Maintenance shutdown (January 2016)
  • Date of next meeting : January 5th

Minutes

Present: Craig, Chris R, Paolo, Ryan, Victoria, Colin,

Review :

  • Colin hopes that the vacuum work for North side is finished.
  • Absorber group are planning to meet tomorrow morning to schedule their activities
  • Absorber filled with Xe on Friday. All is stable and we have been taking data smoothly since Sunday.

Upcoming activities :

Next week's work plan :

  • Activities post-Thursday will probably involve preparation to take FC2 out to get at the absorber to fix it.
  • Beam ends on Friday @ 8 am. Data-taking will stop at 10 pm on the previous evening.
  • We decided to remove the Xe to get a reference data set with an empty absorber. This will happen on Thursday morning.
  • Test of Q4 flush and Temperature readout for Q5 / Q6 will be managed by Henry on Friday 18th.
  • Pierrick requests half a day with the FC at low current. Will discuss with Colin and the absorber team tomorrow.
  • Network configuration for C&M : early to mid January but needs dates to lock it down.
    • Action: ask Pierrick to produce a document on what changes are needed in the network configuration so that system owners
      understand the impact. This should be emailed to system owners ASAP.

Christmas shutdown plans : what needs to be done by Friday and what to monitor over Christmas (Victoria)

  • We have designated end of Wednesday 23rd as "end of play"
  • Victoria will organise the shutdown checklist by looking at the old MOM walkabout list and then considering the new systems.
    Amongst other things that need monitoring over Christmas are
    • Helium levels and pressure in solenoids
    • Focus coil state
    • Cerenkov gas flow
    • Water in quads

January plans

  • A discussion of plans shown in the To-do list was discussed. Timeline and action items can be found
    on that page.

Updated by Boyd, Steven almost 8 years ago ยท 21 revisions